# The DNA of nuclear models: How AI predicts nuclear masses

### v1

@@ Abstract
Obtaining high-precision predictions of nuclear masses, or equivalently nuclear binding energies, $E_{b}$, remains an important goal in nuclear-physics research. Recently, many AI-based tools have shown promising results on this task, some achieving precision that surpasses the best physics models. However, the utility of these AI models remains in question given that predictions are only useful where measurements do not exist, which inherently requires extrapolation away from the training (and testing) samples. Since AI models are largely black boxes, the reliability of such an extrapolation is difficult to assess. We present an AI model that not only achieves cutting-edge precision for $E_{b}$, but does so in an interpretable manner. For example, we find (and explain why) that the most important dimensions of its internal representation form a double helix, where the analog of the hydrogen bonds in DNA here link the number of protons and neutrons found in the most stable nucleus of each isotopic chain. Furthermore, we show that the AI prediction of $E_{b}$ can be factorized and ordered hierarchically, with the most important terms corresponding to well-known symbolic models (such as the famous liquid drop). Remarkably, the improvement of the AI model over symbolic ones can almost entirely be attributed to an observation made by Jaffe in 1969. The end result is a fully interpretable data-driven model of nuclear masses.
@@
In this Letter, we present an AI model—a slightly simplified version of the model from Refs. [23, 24]—that not only achieves cutting-edge precision for predicting nuclear masses, but does so in an interpretable manner. For example, we find (and explain why) that the most important dimensions of its internal representation form a double helix (see Fig. 1), where the analog of the hydrogen bonds in DNA here link the number of protons and neutrons found in the most stable nucleus of each isotopic chain. Furthermore, we show that the AI prediction of $E_{b}$ can be factorized and ordered hierarchically, with the most important terms corresponding to well-known symbolic models, such as the famous liquid drop (LD) [31, 32]. Remarkably, we find that the improvement of the AI model over symbolic ones can almost entirely be attributed to an observation made by Jaffe and collaborators in 1969 [33]. [^2] The relevant sections in Ref. [33] were the outcome of Jaffe’s Junior Paper at Princeton (private communication). The end result is an interpretable data-driven model of nuclear masses.
@@
$-\alpha_{a}\frac{(N-Z)^{2}}{A}+\alpha_{p}\delta(Z,N)\,$, (1)
@@
Remarkably, this same factorization observation was made in 1969 [33], though it seems to have been forgotten until rediscovered by our AI model. Ref. [33] was based on the Garvey-Kelson (GK) relations, patterns found in nuclear data where combinations of $E_{b}$ from a set of nuclei approximately cancel almost anywhere in the nuclear plane (examples in [35]). The GK relations arise because single-nucleon energy levels do not vary much in a small region of the nuclear plane. Any contribution to $E_{b}$ that varies slowly over the nuclear plane, e.g., macroscopic terms like Eq. (The DNA of nuclear models: How AI predicts nuclear masses), approximately cancels in the GK relations. However, microscopic terms that exactly satisfy the full set of GK relations studied in Ref. [33] are severely restricted, they must factorize as in Eq. (5). We henceforth refer to this as Jaffe factorization.
@@
Local Jaffe corrections using the ++ nearest and next-to-nearest neighbors as in Fig. 3 (right) reduces the RMS of our simple model, which recall has no physics microscopic terms, to 0.17 MeV. This is substantially better than WS4, the best physics model, and even slightly better than Ref. [40]. Instead, applying these corrections to WS4 reduces its RMS to 0.16 MeV, which beats all but the most advanced AI models (such as Ref. [23]) without using any AI or complicated interpolation.
Local Jaffe corrections can be improved by also correcting the neutronic (protonic) component for isotopic (isotonic) neighbors, which are used to obtain the protonic (neutronic) corrections in Fig. 3 (right). This is done by considering all available nearest and next-to-nearest neighbors of nucleus $(Z,N)$ in a linear regression that determines local Jaffe corrections for each $Z\in[Z-2,Z+2]$ and $N\in[N-2,N+2]$ [35]. [^6] Local Jaffe corrections work well even when most neighboring nuclei are unobserved, which is a major advantage over using the GK relations directly. Applying these corrections to our symbolic model gives an RMS of 0.14 MeV. Instead, applying them to WS4 produces an RMS of 0.12 MeV (residuals shown in Fig. 4). These are state-of-the-art results for $E_{b}$ predictions—and they are based on explainable corrections to a physics model, not a black-box AI approach. Therefore, our main result is that AI has rediscovered a long forgotten nuclear property we refer to as Jaffe factorization, and when this is applied locally to physics-model calculations it provides state-of-the-art, interpretable predictions for nuclear masses.
@@
The GK relations are the type of empirically derived patterns from data that an AI model is expected to learn, and the (local) Jaffe factorization solution is also something such models are good at discovering. However, without the use of an architecture like ours—chosen to aid interpretability—it would have been difficult to realize what the AI had learned. Finally, given that we now know how optimal local corrections work in simple physics-based terms, it may be possible for nuclear theorists to better understand to what level such corrections can be trusted for highly unstable nuclei.
Acknowledgments—We are grateful to Bob Jaffe for informing us about his undergrad work and pointing us to the factorization presented in Ref. [33]. This work was supported by NSF grant PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). ST was supported by the Office of High Energy Physics of the US Department of Energy (DOE) under Grant No. DE-SC0012567, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (FNAL 20-17). ST is additionally supported by the Swiss National Science Foundation project number PZ00P2_223581, and acknowledges CERN TH Department for hospitality while this research was being carried out.
@@ IV.1 Setting the Scale in the Representation Space
IV.2 Forming Spirals
The same reasoning that prevents the first PC from collapsing under the regularization pressure also applies to the second and third PC dimensions shown in Fig. 1; we will show this later in this section. However, this does not explain the origins of the spiral structure. Naively, there are no cyclic or oscillatory terms in Eq. (The DNA of nuclear models: How AI predicts nuclear masses), making it surprising that the second and third most important PCs would form an oscillatory structure.
To understand why an oscillatory structure is optimal here, we next consider the second most important term in Eq. (The DNA of nuclear models: How AI predicts nuclear masses), the asymmetry term, which contains a factor of $(Z-N)^{2}$. This term is invariant under transformations of the form $Z\to Z+\gamma,N\to N+\gamma$, which correspond to translations along the $\hat{\mathcal{Z}}_{1}$ and $\hat{\mathcal{N}}_{1}$ axes (which are nearly aligned) in Fig. 1. What matters when computing $(Z-N)^{2}$ is only the relative difference in $Z$ and $N$; thus, the same solution can be repeated all along the first PC axis, which is desirable due to the regularization pressure. In principle our AI model could learn to store $(Z-N)^{2}$ for each $(Z,N)$ pair, but the regularization pressure forces it to instead find a more efficient solution, and the spiral is the most efficient option.
@@ V Physics Origins of Jaffe Factorization
Figure S5: Examples of GK relations in the $(Z,N)$ plane. The relative sign (addition or subtraction) of the contribution of each $E_{b}$ value is labeled on the panel (also colored red for addition, blue for subtraction).
@@ VII Jaffe Factorization & Local Corrections
Finally, we find a small improvement by only using neighbors on the same side of the $Z=N$ line as the target nucleus when deriving the local Jaffe corrections. In addition, for nuclei with $N=Z$ only using their nearest neighbors also with $N=Z$ and the same parity (even-even or odd-odd) seems to work the best. These small modifications are included in Fig. S7, though the impact on the RMS is only a reduction of about 5 keV and falls below the level of rounding presented in the main text.
@@ VIII Improving on Basic Local Jaffe Corrections
Applying this proceduce—except for nuclei on or within one unit from the $Z=N$ line, and nuclei with less than 9 measured neighbors, where we instead apply the basic Jaffe procedure from the previous section—to the WS4 predictions reduces the RMS down to 120 keV. We are unaware of any unbiased RMS values that are this low, including from advanced AI models. (Note that some AI works include training-data nuclei in their RMS numbers, which can be highly biased.) If we instead apply this to our base model, the RMS is 144 keV, which given this model has no physics micro-corrections is an incredible result. The residuals for both WS4 and our simple symbolic model with these local Jaffe corrections applied are shown in Fig. S8.
Figure S8: Residuals from (left) our symbolic model and (right) the WS4 model both with local Jaffe corrections applied. Note that the color scale is linear on [1,1][-1,1] and logarithmic otherwise.
In principle, it may be possible to do even better by considering a wider range of neighboring nuclei and allowing some form of variation to the energy-level corrections. However, we note that this RMS is getting close to the value to which the GK relations themselves actually hold. For example, considering only nuclei that have all nearest and next-to-nearest neighbors observed, the local Jaffe procedure here applied to the WS4 model gives an RMS of 78 keV, whereas the GK procedure itself, using a mean of all 12 GK-type predictions for each nucleus [54], has an RMS of 86 keV on these same nuclei. The real advantage of our procedure over using the GK relations directly is that local Jaffe corrections also work well when most of the neighboring nuclei are not measured. As noted in the Letter, scientifically what is desired is not higher-precision predictions of nuclei surrounded by well-measured neighbors, but trustworthy ways of extrapolating corrections away from the stable region. The local Jaffe approach provides an interpretable method for deriving such corrections.

### v2

@@ Abstract
Obtaining high-precision predictions of nuclear masses, or equivalently nuclear binding energies, $E_{b}$, remains an important goal in nuclear-physics research. Recently, many AI-based tools have shown promising results on this task, some achieving precision that surpasses the best physics models. However, the utility of these AI models remains in question given that predictions are only useful where measurements do not exist, which inherently requires extrapolation away from the training (and testing) samples. Since AI models are largely black boxes, the reliability of such an extrapolation is difficult to assess. We present an AI model that not only achieves cutting-edge precision for $E_{b}$, but does so in an interpretable manner. For example, we find that (and explain why) the most important dimensions of its internal representation form a double helix, where the analog of the hydrogen bonds in DNA here link the number of protons and neutrons found in the most stable nucleus of each isotopic chain. Furthermore, we show that the AI prediction of $E_{b}$ can be factorized and ordered hierarchically, with the most important terms corresponding to well-known symbolic models (such as the famous liquid drop). Remarkably, the improvement of the AI model over symbolic ones can almost entirely be attributed to an observation made by Jaffe in 1969 based on the structure of most known nuclear ground states. The end result is a fully interpretable data-driven model of nuclear masses based on physics deduced by AI.
@@
In this Letter, we present an AI model—a slightly simplified version of the model from Refs. [23, 24]—that not only achieves cutting-edge precision for predicting nuclear masses, but does so in an interpretable manner. For example, we find (and explain why) that the most important dimensions of its internal representation form a double helix (see Fig. 1), where the analog of the hydrogen bonds in DNA here link the number of protons and neutrons found in the most stable nucleus of each isotopic chain. Furthermore, we show that the AI prediction of $E_{b}$ can be factorized and ordered hierarchically, with the most important terms corresponding to well-known symbolic models, such as the famous liquid drop (LD) [31, 32]. Remarkably, we find that the improvement of the AI model over symbolic ones can almost entirely be attributed to an observation made by Jaffe and collaborators in 1969 [33]. [^2] The relevant sections in Ref. [33] were the outcome of Jaffe’s Junior Paper at Princeton (private communication). The end result is a model of nuclear masses based on physics deduced by AI.
@@
$-\alpha_{a}\frac{(N-Z)^{2}}{A}+\alpha_{p}\frac{\delta(Z,N)}{A^{1/2}}\,$, (1)
@@
Remarkably, this same factorization observation was made in 1969 [33], though it seems to have been forgotten until rediscovered by our AI model. Ref. [33] was based on the Garvey-Kelson (GK) relations, patterns found in nuclear data where combinations of $E_{b}$ from a set of nuclei approximately cancel almost anywhere in the nuclear plane (examples in [35]). The GK relations arise because single-nucleon energy levels do not typically vary much in a small region of the nuclear plane. Any contribution to $E_{b}$ that varies slowly over the nuclear plane, e.g., macroscopic terms like Eq. (The DNA of nuclear models: How AI predicts nuclear masses), approximately cancels in the GK relations. However, microscopic terms that exactly satisfy the full set of GK relations studied in Ref. [33] are severely restricted, they must factorize as in Eq. (5). We henceforth refer to this as Jaffe factorization.
@@
Local Jaffe corrections using the ++ nearest and next-to-nearest neighbors as in Fig. 3 (right) reduces the RMS of our simple model, which recall has no physics microscopic terms, to 0.18 MeV. This is substantially better than WS4, the best physics model, and even consistent with Ref. [40]. Instead, applying these corrections to WS4 reduces its RMS to 0.16 MeV, which beats all but the most advanced AI models (such as Ref. [23]) without using any AI or complicated interpolation.
Local Jaffe corrections can be improved by also correcting the neutronic (protonic) component for isotopic (isotonic) neighbors, which are used to obtain the protonic (neutronic) corrections in Fig. 3 (right). This is done by considering all available nearest and next-to-nearest neighbors of nucleus $(Z,N)$ in a linear regression that determines local Jaffe corrections for each $Z\in[Z-2,Z+2]$ and $N\in[N-2,N+2]$ [35]. [^6] Local Jaffe corrections work well even when most neighboring nuclei are unobserved, which is a major advantage over using the GK relations directly; see Table S1. Applying these corrections to WS4 produces an RMS below 0.12 MeV (residuals shown in Fig. 4). These are state-of-the-art results for $E_{b}$ predictions—and they are based on explainable corrections to a physics model, not a black-box AI approach. Therefore, our main result is that AI has rediscovered a long forgotten nuclear property we refer to as Jaffe factorization, and when this is applied locally to physics-model calculations it provides state-of-the-art, interpretable predictions for nuclear masses.
@@
The GK relations are the type of empirically derived patterns from data that an AI model is expected to learn, and the (local) Jaffe factorization solution is also something such models are good at discovering. However, without the use of an architecture like ours—chosen to aid interpretability—it would have been difficult to realize what the AI had learned. Given that we now know how optimal local corrections work in simple physics-based terms, it should be possible for nuclear theorists to better predict the highly unstable nuclei for which such corrections can be trusted. Our work demonstrates how human-AI collaboration can unlock new scientific understanding and accelerate discovery.
Acknowledgments—We are grateful to Bob Jaffe for informing us about his undergrad work, pointing us to the factorization presented in Ref. [33], and providing comments on our work. We also thank Jesse Thaler, Antoine Belley, and Jose Munoz for providing useful feedback. This work was supported by NSF grant PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). S.T. was supported by the Office of High Energy Physics of the US Department of Energy (DOE) under Grant No. DE-SC0012567, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (FNAL 20-17). S.T. is additionally supported by the Swiss National Science Foundation project number PZ00P2_223581, and acknowledges CERN TH Department for hospitality while this research was being carried out.
@@ IV.1 Setting the Scale in the Representation Space
IV.2 Forming Helices
The same reasoning that prevents the first PC from collapsing under the regularization pressure also applies to the second and third PC dimensions shown in Fig. 1; we will show this later in this section. However, this does not explain the origins of the helix structure. Naively, there are no cyclic or oscillatory terms in Eq. (The DNA of nuclear models: How AI predicts nuclear masses), making it surprising that the second and third most important PCs would form an oscillatory structure.
To understand why an oscillatory structure is optimal here, we next consider the second most important term in Eq. (The DNA of nuclear models: How AI predicts nuclear masses), the asymmetry term, which contains a factor of $(Z-N)^{2}$. This term is invariant under transformations of the form $Z\to Z+\gamma,N\to N+\gamma$, which correspond to translations along the $\hat{\mathcal{Z}}_{1}$ and $\hat{\mathcal{N}}_{1}$ axes (which are nearly aligned) in Fig. 1. What matters when computing $(Z-N)^{2}$ is only the relative difference in $Z$ and $N$; thus, the same solution can be repeated all along the first PC axis, which is desirable due to the regularization pressure. In principle our AI model could learn to store $(Z-N)^{2}$ for each $(Z,N)$ pair, but the regularization pressure forces it to instead find a more efficient solution, and the helix is the most efficient option.
@@ V Physics Origins of Jaffe Factorization
Figure S5: Examples of GK relations in the $(Z,N)$ plane for Eqs. (S24) and (S25). The relative sign (addition or subtraction) of the contribution of each $E_{b}$ value is labeled on the panel (also colored red for addition, blue for subtraction).
@@ VII Jaffe Factorization & Local Corrections
Finally, we find a small improvement by not using odd-odd $Z=N$ neighbors when deriving the local Jaffe corrections. In addition, we do not attempt to correct odd-odd $Z=N$ nuclei. These improvements, e.g.,reduce the RMS from 160 to 157 keV for the (next-to-)nearest isoto(p,n)ic neighbors. These small modifications are included in Fig. S7, though the impact on the RMS is small.
@@ VIII Improving on Basic Local Jaffe Corrections
Applying this procedure—except for $Z=N$ odd-odd nuclei, and nuclei with less than 7 total and 3 isoto(p,n)ic measured neighbors (of the 24 possible), where we instead apply the basic Jaffe procedure from the previous section—to the WS4 predictions reduces the RMS down to 115 keV. The residuals are shown in Fig. 4. We are unaware of any unbiased RMS values that are this low, including from advanced AI models. (Note that some AI works include training-data nuclei in their RMS numbers, which can be highly biased.)
Table S1 shows how our procedure compares with using the GK relations directly to correct WS4. For nuclei that have nearly all nearest and next-to-nearest neighbors observed, the local Jaffe procedure is similar to GK using a mean of all 12 GK-type predictions for each nucleus [54]. The real advantage of our procedure over using the GK relations directly is that local Jaffe corrections also work well when most of the neighboring nuclei are not measured, which is the most interesting region. As noted in the Letter, scientifically what is desired is not higher-precision predictions of nuclei surrounded by well-measured neighbors, but trustworthy ways of extrapolating corrections away from the stable region. The local Jaffe approach provides an interpretable method for deriving such corrections.
Number (Next-to-)Nearest Neighbors WS4 RMS GK Corrected RMS Local Jaffe Corrected RMS
20–24 268 keV 93 keV 94 keV
11–19 300 keV 161 keV 160 keV
1–10 391 keV 293 keV 207 keV
Table S1: Performance of the WS4 model without (middle) and with (right) local Jaffe corrections for nuclei categorized by how many of their (next-to-)nearest neighbors are included in the data sample. If all of the (next-to-)nearest neighbors are observed, the value is 24. As expected the performance of the WS4 model, along with both GK and our local Jaffe corrections degrades when moving towards the edges of the observed nuclear plane; however, even when most of the neighbors are not observed, the local Jaffe procedure still drastically improves the precision of the nuclear mass predictions. In addition, our Local Jaffe corrections outperform the GK-based corrections for nuclei with few observed neighbors.